Mrs. Morris has 84 dance students. If she has five times as many teenage students as adult students, how many adult students does she have?
Answer by edjones(7311)

a=adult t=teens
t=5a
t+a=84
substitute 5a for t in the 2nd equation: 5a+a=84
6a=84
divide 6 into both sides: a=84/6=14 adults
Check:
14*5=70 teens
70+14=84
Ed